OpenStudy (anonymous):

What is the length of the leg in a right triangle whose hypotenuse is 13 and other leg is 12. A. 3 B. 5 C. 4 D. 2

OpenStudy (anonymous):

@AdamK

OpenStudy (kc_kennylau):

(one leg)^2 + (another leg)^2 = (hypotenuse)^2

OpenStudy (anonymous):

@moonmonsterz my feels

OpenStudy (kc_kennylau):

So we have (the leg)^2 + 12^2 = 13^2

OpenStudy (adamk):

Didn't we just do this one?

OpenStudy (kc_kennylau):

I hope you can continue the calculation yourself xd

OpenStudy (adamk):

What's 169 - 144?

OpenStudy (anonymous):

@AdamK no there similar questions

OpenStudy (adamk):

Now take the square root of that and you have your answer

OpenStudy (kc_kennylau):

Is no admin here @ganeshie8